1) A solid ball of radius 7.25 cm and mass 7.43kg starts from rest and rolls without slipping down a 20.7 degree incline that is 1.22m long. Calculate the angular velocity (in rad/s) of the ball when it reaches the bottom
2:43 PM o 37 BD ooooo AT&T LTE 2)A uniform thin ro
Answer #1

Q1.

mass=m=7.43 kg

radius=r=7.25 cm=0.0725 m

let final speed at the bottom be v m/s.

then final angular speed=v/r rad/s

moment of inertia of a solid sphere=0.4*mass*radius^2

balancing total mechanical energy:

initial potential energy=final linear kinetic energy + final rotational kinetic energy

==>mass*g*height=0.5*mass*v^2+0.5*moment of inertia*angular speed^2

==>mass*g*height=0.5*mass*v^2+0.5*0.4*mass*r^2*(v/r)^2

==>9.8*1.22*sin(20.7)=0.5*v^2+0.2*v^2=0.7*v^2

==>v=sqrt(9.8*1.22*sin(20.7)/0.7)=4.0517 m/s